SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (August 8, 2006)

Attendees: Rainer Bartoldus, Al Eisner, Arik Kreisel, Selina Li, Dong Su, Mark Tibbetts, Dan Walker

Operations

  • L1 trigger rate dissapeared on the FCTS disaply on Monday (August 7) around 5am. This is at least the second or third times of occurance. Su Dong suggested the shifter to page ODF instead and reconfiguring ORC solved the problem. This was a problem with monitoring and the data were fine. This problem has been under investigation by the ODF group.
  • The ZPD Tandip refernce plot went negative in the y-axis due to a JAS overlaid problem (see plot). However, this happened only briefly in a run, and the plot appeared to be normal again after the JAS page was refreshed a few times.

EMT

  • Dan reported that the deadtime contribution from EMT basically has shown no significant impact after changing the prescale from 3 to 1 last Monday (July 31). We will leave the prescale to be 1 from now on unless it is necessary to change in the future in which case we could easily change back to use a prescale of 3.

Online

  • Jim Hamilton installed the new online release at IR2 this morning. The 18.7.0c built is finally in!

L3

  • This new online release (based on the 18.7.0c built) doesn't include the L3 tracking derivative changes since the number of reco tracks went down and still needed to be fully understood.
  • We had a discussion of the ghost tracks found by L3. Rainer pointed out that L3 picks up only about 50% of recon tracks that are ghost tracks. Looking more into this, it seems like it's a problem with linking track not ghost tracks so unlikely a problem with L3 tracking.

Data Quality and Monitoring

  • Rahmat was unable to make it to the meeting today, but he uploaded the lumi ratio plots for us to look at. The plots look reasonable, except that we would prefer the x-axis to be in terms of time or date. The lables for BC mupair should be specified in all 4 plots.

Trigger Configuration

  • The proposed trigger configuration change involves a change in the ZPD configuration. Rainer is contacting the relevant people to get more info about how to make the ZPD configuration change and then load into the configuration database. Once we take data with the new trg configuration, we would also need the corresponding MC simulations.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.